OpenStudy (anonymous):

3.Solve x2 + 18 = –9x. (1 point) 4.Solve 4x2 = 12x please help by solving completing the square please show work

OpenStudy (anonymous):

x^2+9x+18=0 x^2+6x+3x+18=0 solve further by yourself 4. 4x^2-12x=0 x(4x-12)=0 x=0 , further solve by yourself